For a car loan using simple interest at a given rate, the amount of interest charged varies jointly with the loan amount (also known as the principal) and the time of the loan (in years). If a $15,000 car loan earns $2,610 in simple interest over 6 years, how much interest will $32,000 car loan earn over 7 years?
Answer by stanbon(75887)

For a car loan using simple interest at a given rate, the amount of interest charged varies jointly with the loan amount (also known as the principal) and the time of the loan (in years). If a $15,000 car loan earns $2,610 in simple interest over 6 years, how much interest will $32,000 car loan earn over 7 years?
-----------------
I = P*r*t
----
Solve for "r":
2610 = 15000*r*6
r = 2610/90000 = 0.029
------
Solve for "I"
I = 32000*0.029*7
I = $6496
===================
Cheers,
Stan H.------------